OpenStudy (anonymous):

Let f(x)=5·x2−2·x. Find f(x+h)−f(x)/h if h≠0

OpenStudy (campbell_st):

replace x with (x+h) in every term containing an x f(x + h) = 5(x+h)^2 - 2(x+h) expand and simplify 5(x^2 +2xh + h^2) -2x -2h = 5x^2 + 10xh + 5h^2 -2x -2h Now f(x+h) -f(x) = 5x^2 + 10xh + 5h^2 - 2x - 2h - 5x^2 - 2x leaves 10xh +5h^2 -2h divide by h \[(10xh - 5h^2 -2h)/h = 10x - 5h - 2\] this is differentiation from 1st principles...

OpenStudy (anonymous):

Thanks, I forgot to substitute one of the x +h. I appreciate the help